# A topological constraint on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v1

@@ 3.1 General case
We will combine Proposition 2.8 with the results in [4, 15, 16] about the knot DGA. Some conventions about the knot DGA in these papers are slightly different from each other, as summarized in [16, Appendix A]. In this section, we mainly refer to [16] for the definition of the knot DGA.
Let $K$ be an oriented knot in $\mathbb{R}^{3}$ given by the closure of a braid $B$. Then, a DGA over $R$, called the knot DGA of $K$, is constructed from $B$ as in [16, Definition 3.11]. Its stable tame isomorphism class as a DGA over $R$ depends only on the isotopy class of $K$ as an oriented knot. See [15, Theorem 1.4] and [16, Theorem 3.1]. Let us denote this DGA by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ and its homology $\operatorname{Ker}\partial_{K}/\operatorname{Im}\partial_{K}$ by $HC_{*}(K)$. We will review a description of the 00-th degree part $HC_{0}(K)$ in Section 3.2.
@@
The base change by $\psi^{\prime}$ is due to the difference of conventions between [4] and [16]. See [16, Appendix A]. It follows from this theorem that there are unital graded ring homomorphisms
@@
We continue to consider the setup in Subsection 2.5.2. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$. For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. Then we have the following theorem analogous to [17, Proposition 4.6].
@@
for every $x\in\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})$ and $r\in R$, where $\psi$ a ring isomorphism defined by
@@
A unital graded ring homomorphism $\Psi_{\varphi}$ satisfying $\Psi_{\varphi}\circ\partial_{K_{0}}=\partial_{K_{1}}\circ\Psi_{\varphi}$ is defined as the composition of the following morphisms of differential graded rings:
[…]
@@
and $\psi$ is described as in the theorem for the integers $a,b,c$ and $d$ which determine (3). ∎
We recall the notion of an augmentation of a DGA. See, for instance, [16, Definition 2.6].
@@
Let $\mathbb{k}$ be a field and $(\mathcal{A}_{*},\partial)$ be a DGA over $R$. A ring homomorphism $\varepsilon\colon\mathcal{A}_{*}\to\mathbb{k}$ is called an augmentation of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ if $\varepsilon(\mathbb{1})=1$, $\varepsilon(x)=0$ for any $x\in\mathcal{A}_{p}$ with $p\neq 0$, and $\varepsilon\circ\partial\colon\mathcal{A}_{1}\to\mathbb{k}$ is the zero map. Here, $\mathbb{1}$ is the unit of $\mathcal{A}_{*}$.
A knot invariant is defined from the set of augmentations of the knot DGA to $\mathbb{C}$.
@@
$V_{K}\coloneqq\left\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\ \middle|\begin{array}[]{l}\text{There exists an augmentation }\varepsilon\text{ of }(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})\text{ to }\mathbb{C}\\ \text{such that }\varepsilon(\lambda)=x,\ \varepsilon(\mu)=y\text{ and }\varepsilon(U)=Z\end{array}\right\}.$
Here, $\varepsilon(r\cdot\mathbb{1})$ for any $r\in R$ is abbreviated by $\varepsilon(r)$.
From Theorem 3.2, we obtain a refinement of [17, Theorem 1.1].
@@
There exist integers $b^{\prime},c^{\prime}$ and $d^{\prime}$ for which we have an injective map
$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(xy^{b^{\prime}}Z^{c^{\prime}},yZ^{d^{\prime}},Z).$
@@
By Theorem 3.2, for any augmentation $\varepsilon$ of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{1}),\partial_{K_{1}})$ to $\mathbb{C}$ with $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$, the composite map $\varepsilon\circ\Psi_{\varphi}\colon\mathcal{A}_{*}^{\mathrm{knot}}(K_{0})\to\mathbb{C}$ is an augmentation of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ to $\mathbb{C}$ such that
@@
From the definition of $V_{K_{i}}$ for $i\in\{0,1\}$, this implies that we have a map
$V_{K_{1}}\to V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
We use the following fact by [15, Proposition 4.3]: Up to chain isomorphisms of DGA over $R$, $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ is unchanged by the substitution $\lambda\mapsto\lambda^{-1}U$, $\mu\mapsto\mu^{-1}U$ and $U\mapsto U$. Hence, we have a bijection
$V_{K_{0}}\to V_{K_{0}}\colon(x,y,Z)\mapsto(x^{-1}Z,y^{-1}Z,Z).$
If $a=1$, we take $b^{\prime}\coloneqq b$, $c^{\prime}\coloneqq c$ and $d^{\prime}\coloneqq d$. If $a=-1$, we take $b^{\prime}\coloneqq-b$, $c^{\prime}\coloneqq-c+1$ and $d^{\prime}\coloneqq-d+1$. Then, the assertion of the corollary holds. ∎
3.2 Proof of Theorem 1.1
First, let us review how to compute $HC_{0}(K)$ for an oriented knot $K$. We refer to [16, Section 3]. Note that since $\mathcal{A}^{\mathrm{knot}}_{p}(K)=0$ for $p<0$, $HC_{0}(K)=\mathcal{A}^{\mathrm{knot}}_{0}(K)/\partial_{K}(\mathcal{A}^{\mathrm{knot}}_{1}(K))$.
For any $n\in\mathbb{Z}_{\geq 1}$, let $B_{n}$ denote the group of braids on $n$ strands, which is generated by $\{\sigma_{1},\dots,\sigma_{n-1}\}$ with relations $\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1}$ for $21\leq i\leq n-$ 2 and $\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}$ for $|i-j|\geq 2$. Let $\mathcal{A}_{n}$ be the $R$-algebra freely generated by the set $\{a_{i,j}\mid i,j\in\{1,\dots,n\},\ i\neq j\}$. Then, a group homomorphism $B_{n}\to\mathrm{Aut}(\mathcal{A}_{n})\colon B\mapsto\phi_{B}$ is defined as in [16, Definition 3.3]. It is determined by $\phi_{\sigma_{k}}\colon\mathcal{A}_{n}\to\mathcal{A}_{n}$ for each generator $\sigma_{k}$, and $\phi_{\sigma_{k}}$ is given by
$\phi_{\sigma_{k}}\colon\begin{cases}a_{i,j}\mapsto a_{i,j}&\text{if }i,j\neq k,k+1,\\ a_{k+1,i}\mapsto a_{k,i}&\text{ if }i\neq k,k+1,\\ a_{i,k+1}\mapsto a_{i,k}&\text{ if }i\neq k,k+1,\\ a_{k,k+1}\mapsto-a_{k+1,k},&\\ a_{k+1,k}\mapsto-a_{k,k+1},&\\ a_{k,i}\mapsto a_{k+1,i}-a_{k+1,k}a_{k,i}&\text{ if }i\neq k,k+1,\\ a_{i,k}\mapsto a_{i,k+1}-a_{i,k}a_{k,k+1}&\text{ if }i\neq k,k+1.\end{cases}$ (10)
See also [7, Section 2.1]. We remark that $\phi_{B\cdot B^{\prime}}=\phi_{B}\circ\phi_{B^{\prime}}$ for any $B,B^{\prime}\in B_{n}$.
The automorphism $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n})$ for $B\in B_{n}$ is extended to an automorphism on $\mathcal{A}_{n+1}$ as follows: We consider the group homomorphism $B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$ defined as above and an injective map $B_{n}\to B_{n+1}$ defined by adding the extra $(n+1)$-th strand to any braid on $n$ strands (see [16, Remark 3.5]). Then, for any $B\in B_{n}$, $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n+1})$ is defined via the composite map $B_{n}\hookrightarrow B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$. In addition, let us label the $(n+1)$-st strand by * and denote $a_{i,n+1}$ by $a_{i,*}$ and $a_{n+1,i}$ by $a_{*,i}$ for every $i\in\{1,\dots,n\}$.
@@
$\begin{array}[]{ll}\mathbb{A}_{i,j}\coloneqq\begin{cases}a_{i,j}&\text{ if }i<j,\\ 1-\mu&\text{ if }i=j,\\ -\mu a_{i,j}&\text{ if }i>j,\end{cases}&\widehat{\mathbb{A}}_{i,j}\coloneqq\begin{cases}Ua_{i,j}&\text{ if }i<j,\\ U-\mu&\text{ if }i=j,\\ -\mu a_{i,j}&\text{ if }i>j.\end{cases}\end{array}$
Fix $B\in B_{n}$ which has the writhe $w$ (the sum of the exponents in the braid word). Let $\bm{\Lambda}$ be an $n\times n$ diagonal matrix with $\bm{\Lambda}_{1,1}=\lambda\mu^{w}U^{-(w-n+1)/2}$ and $\bm{\Lambda}_{i,i}=1$ for every $i\in\{2,\dots,n\}$. Let $\bm{\Phi}^{L}_{B}$ be an $n\times n$ matrix whose entries $(\bm{\Phi}^{L}_{B})_{i,j}\in\mathcal{A}_{n}$ for $i,j\in\{1,\dots,n\}$ are determined by
@@
Let $K$ be the oriented knot in $\mathbb{R}^{3}$ given by the closure of the braid $B$. By [16, Theorem 4.1, Remark 4.2], $HC_{0}(K)$ is the quotient of $\mathcal{A}_{n}$ by the two-sided ideal generated by all entries of the two matrices
@@
where $(\phi_{B}(\mathbb{A}))_{i,j}\coloneqq\phi_{B}(\mathbb{A}_{i,j})$ for $i,j\in\{1,\dots,n\}$. The isomorphism class of $HC_{0}(K)$ as a unital $R$-algebra depends only on the isotopy class of $K$. See [16, Theorem 3.15] and [15, Theorem 1.4 (2)].
We consider a commutative $R$-algebra $HC_{0}^{\mathrm{ab}}(K)$ defined as the quotient of $HC_{0}(K)$ by the relations $\xi\eta=\eta\xi$ for all $\xi,\eta\in HC_{0}(K)$. Using the $R$-algebra $R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]$ of polynomials and a natural map $\pi_{n}\colon\mathcal{A}_{n}\to R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]\colon a_{i,j}\mapsto a_{i,j}$, we have

### v2

@@ 3.1 General case
We will combine Proposition 2.8 with the results in [3, 14, 15] about the knot DGA. Some conventions about the knot DGA in these papers are slightly different from each other, as summarized in [15, Appendix A]. In this section, we mainly refer to [15] for the definition of the knot DGA.
Let $K$ be an oriented knot in $\mathbb{R}^{3}$ given by the closure of a braid $B$. Then, a DGA over $R$, called the knot DGA of $K$, is constructed from $B$ as in [15, Definition 3.11]. Its stable tame isomorphism class as a DGA over $R$ depends only on the isotopy class of $K$ as an oriented knot. See [14, Theorem 1.4] and [15, Theorem 3.1]. Let us denote this DGA by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ and its homology $\operatorname{Ker}\partial_{K}/\operatorname{Im}\partial_{K}$ by $HC_{*}(K)$. We will review a description of the 0-th degree part $HC_{0}(K)$ in Section 3.2.
@@
The base change by $\psi^{\prime}$ is due to the difference of conventions between [3] and [15]. See [15, Appendix A]. It follows from this theorem that there are unital graded ring homomorphisms
@@
We continue to consider the setup in Subsection 2.5.2. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$. For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. Then we have the following theorem analogous to [16, Proposition 4.6].
@@
for every $x\in\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})$ and $r\in R$, where $\psi$ is a ring isomorphism defined by
@@
A unital graded ring homomorphism $\Psi_{\varphi}$ satisfying $\Psi_{\varphi}\circ\partial_{K_{0}}=\partial_{K_{1}}\circ\Psi_{\varphi}$ is defined by the composition of the following morphisms of differential graded rings:
${(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\bar{I}_{K_{0}}}$${(\mathcal{A}_{*}(\Lambda_{K_{0}}),\partial_{J,\mathfrak{s}^{0}_{K_{0}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{I(\mathfrak{s}^{0}_{K_{0}},\mathfrak{s}_{K_{0}})}$${(\mathcal{A}_{*}(\Lambda_{K_{0}}),\partial_{J,\mathfrak{s}_{K_{0}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{L_{\varphi}}}$${(\mathcal{A}^{\mathrm{knot}}_{*}(K_{1}),\partial_{K_{1}})}$${(\mathcal{A}_{*}(\Lambda_{K_{1}}),\partial_{J,\mathfrak{s}^{0}_{K_{1}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{I_{K_{1}}}$${(\mathcal{A}_{*}(\Lambda_{K_{1}}),\partial_{J,\mathfrak{s}_{K_{1}}}).\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{I(\mathfrak{s}^{0}_{K_{1}},\mathfrak{s}_{K_{1}})^{-1}}$
@@
then $\psi$ is described as in the theorem by using the integers $a,b,c$ and $d$ which determine (3). ∎
We recall the definition of augmentations of a DGA. See, for instance, [15, Definition 2.6].
@@
Let $\mathbb{k}$ be a field and let $(\mathcal{A}_{*},\partial)$ be a unital non-commutative DGA over $R$. A ring homomorphism $\varepsilon\colon\mathcal{A}_{*}\to\mathbb{k}$ is called an augmentation of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ if $\varepsilon(\mathbb{1})=1$, $\varepsilon(x)=0$ for any $x\in\mathcal{A}_{p}$ with $p\neq 0$, and $\varepsilon\circ\partial\colon\mathcal{A}_{1}\to\mathbb{k}$ is the zero map. Here, $\mathbb{1}$ is the unit of $\mathcal{A}_{*}$.
In the following, given an augmentation $\varepsilon$ as in the definition, we abbreviate $\varepsilon(r\cdot\mathbb{1})$ by $\varepsilon(r)$ for any $r\in R$.
Let us prepare a lemma which is necessary to deduce Proposition 3.10 and Proposition 3.17.
Lemma 3.4.
Let $(\mathcal{A}_{*},\partial)$ be a unital non-commutative DGA over $R$ and let $\mathbb{k}$ be a field. Let $H^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial)$ denote the commutative $R$-algebra obtained by the abelianization of the 0-th homology of $(\mathcal{A}_{*},\partial)$. Suppose that there exist a polynomial $P\in R[T]$ and a unital $R$-algebra map
$J\colon R[T]/(P)\to H^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial).$
Then, for any augmentation $\varepsilon$ of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)\in(\mathbb{k}^{*})^{3}$, the polynomial $\left.P\right|_{\lambda=x,\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
Proof.
Any augmentation $\varepsilon$ of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$ induces a unital ring homomorphism $\bar{\varepsilon}\colon HC^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial)\to\mathbb{k}$ such that $(\bar{\varepsilon}(\lambda),\bar{\varepsilon}(\mu),\bar{\varepsilon}(U))=(x,y,Z)$. We obtain a unital ring homomorphism $\bar{\varepsilon}\circ J\colon R[T]/(P)\to\mathbb{k}$, and $(\bar{\varepsilon}\circ J)(T)\in\mathbb{k}$ is a root of $\left.P\right|_{\lambda=x,\mu=y,U=Z}\in\mathbb{k}[T]$. ∎
The next proposition is used to prove Theorem 3.15 and Theorem 3.19 by taking $\mathbb{k}=\mathbb{Q}$.
Proposition 3.5.
Suppose that $K_{1}$ is the unknot. Then, for any $y,Z\in\mathbb{k}^{*}$ satisfying $y\neq Z^{n}$ for every $n\in\mathbb{Z}$, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)$.
Proof.
If $K_{1}$ is the unknot, from the trivial braid on a single strand, we can compute as in [15, Example 3.13] that $\mathcal{A}^{\mathrm{knot}}_{0}(K_{1})=R=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ and $\partial_{K_{1}}(\mathcal{A}_{1}^{\mathrm{knot}}(K_{1}))$ is the ideal generated by
$U-\lambda-\mu+\lambda\mu.$
Let us take $a\in\{\pm 1\}$ and $b,c,d\in\mathbb{Z}$ to be the integers in Theorem 3.2. For any $y,Z\in\mathbb{k}^{*}$ satisfying $y\notin\{Z^{n}\mid n\in\mathbb{Z}\}$, we take an augmentation $\varepsilon_{1}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})\to\mathbb{k}$ determined by
$\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=\displaystyle{\frac{Z-(y\cdot Z^{-d})^{a}}{1-(y\cdot Z^{-d})^{a}}},&\varepsilon_{1}(\mu)=(y\cdot Z^{-d})^{a},&\varepsilon_{1}(U)=Z.\end{array}$
Note that $y\cdot Z^{-d}\neq 1$ and $y\cdot Z^{-d}\neq Z^{a}$ from the condition on $y$ and $Z$.
Composing with the map $\Psi_{\varphi}$ of Theorem 3.2, we obtain $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})\to\mathbb{k}$. It is an augmentation of the knot DGA of $K_{0}$ to $\mathbb{k}$ such that
$\begin{array}[]{cc}\varepsilon(\mu)=\varepsilon_{1}(\mu^{a}U^{d})=y,&\varepsilon(U)=\varepsilon_{1}(U)=Z.\end{array}$
Therefore, $\varepsilon$ satisfies the required condition. ∎
Let us also consider a knot invariant defined from $\mathbb{C}$-valued augmentations of the knot DGA.
@@
[…]
From Theorem 3.2, we obtain a refinement of [16, Theorem 1.1].
@@
There exist $a\in\{\pm 1\}$ and integers $b,c$ and $d^{\prime}$ for which we have an injective map
$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},yZ^{d},Z).$
@@
Using $\Psi_{\varphi}$ of Theorem 3.2, it follows that for any augmentation $\varepsilon$ of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{1}),\partial_{K_{1}})$ to $\mathbb{C}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$, the composite map $\varepsilon\circ\Psi_{\varphi}\colon\mathcal{A}_{*}^{\mathrm{knot}}(K_{0})\to\mathbb{C}$ is an augmentation of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ to $\mathbb{C}$ such that
@@
From the definition of $V_{K_{i}}$ for $i\in\{0,1\}$, we obtain a map from $V_{K_{1}}$ to $V_{K_{0}}$ as in the assertion. ∎
3.2 Proof of Theorem 1.1 for $K^{\prime}\#T_{(2,2m+1)}$ ($m\neq 0,-1$)
3.2.1 Review of the definition of $HC_{0}(K)$
First, let us review how to compute $HC_{0}(K)$ for an oriented knot $K$. We refer to [15, Section 3]. Note that since $\mathcal{A}^{\mathrm{knot}}_{p}(K)=0$ for $p<0$, $HC_{0}(K)=\mathcal{A}^{\mathrm{knot}}_{0}(K)/\partial_{K}(\mathcal{A}^{\mathrm{knot}}_{1}(K))$.
For any $n\in\mathbb{Z}_{\geq 1}$, let $B_{n}$ denote the group of braids on $n$ strands, which is generated by $\{\sigma_{1},\dots,\sigma_{n-1}\}$ with relations $\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1}$ for $1\leq i\leq n-2$ and $\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}$ for $|i-j|\geq 2$. Let $\mathcal{A}_{n}$ be the $R$-algebra freely generated by the set $\{a_{i,j}\mid i,j\in\{1,\dots,n\},\ i\neq j\}$. Then, a group homomorphism $B_{n}\to\mathrm{Aut}(\mathcal{A}_{n})\colon B\mapsto\phi_{B}$ is defined as in [15, Definition 3.3]. It is determined by $\phi_{\sigma_{k}}\colon\mathcal{A}_{n}\to\mathcal{A}_{n}$ for each generator $\sigma_{k}$, and $\phi_{\sigma_{k}}$ is given by
[…]
See also [6, Section 2.1]. We remark that $\phi_{B\cdot B^{\prime}}=\phi_{B}\circ\phi_{B^{\prime}}$ for any $B,B^{\prime}\in B_{n}$.
The automorphism $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n})$ for $B\in B_{n}$ is extended to an automorphism on $\mathcal{A}_{n+1}$ as follows: We consider the group homomorphism $B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$ defined as above and an injective map $B_{n}\to B_{n+1}$ defined by adding the extra $(n+1)$-th strand to any braid on $n$ strands (see [15, Remark 3.5]). Then, for any $B\in B_{n}$, $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n+1})$ is defined via the composite map $B_{n}\hookrightarrow B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$. In addition, let us label the $(n+1)$-th strand by * and denote $a_{i,n+1}$ by $a_{i,*}$ and $a_{n+1,i}$ by $a_{*,i}$ for every $i\in\{1,\dots,n\}$.
@@ 3.2.1 Review of the definition of $HC_{0}(K)$
[…]
Fix $B\in B_{n}$ which has the writhe $w$ (the sum of the exponents in the braid word). Let $\bm{\Lambda}$ be an $n\times n$ diagonal matrix such that $\bm{\Lambda}_{1,1}=\lambda\mu^{w}U^{-(w-n+1)/2}$ and $\bm{\Lambda}_{i,i}=1$ for every $i\in\{2,\dots,n\}$. Let $\bm{\Phi}^{L}_{B}$ be an $n\times n$ matrix whose entries $(\bm{\Phi}^{L}_{B})_{i,j}\in\mathcal{A}_{n}$ for $i,j\in\{1,\dots,n\}$ are determined by
@@
Let $K$ be the oriented knot in $\mathbb{R}^{3}$ given by the closure of the braid $B$. By [15, Theorem 4.1, Remark 4.2], $HC_{0}(K)$ is the quotient of $\mathcal{A}_{n}$ by the two-sided ideal generated by all entries of the two matrices
@@
where $(\phi_{B}(\mathbb{A}))_{i,j}\coloneqq\phi_{B}(\mathbb{A}_{i,j})$ for $i,j\in\{1,\dots,n\}$. The isomorphism class of $HC_{0}(K)$ as a unital $R$-algebra depends only on the isotopy class of $K$ as an oriented knot. See [15, Theorem 3.15] and [14, Theorem 1.4 (2)].
We consider a commutative $R$-algebra $HC_{0}^{\mathrm{ab}}(K)$ defined as the quotient of $HC_{0}(K)$ by the relations $\xi\eta=\eta\xi$ for all $\xi,\eta\in HC_{0}(K)$. Using the $R$-algebra $R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]$ of polynomials and a natural map
$\pi_{n}\colon\mathcal{A}_{n}\to R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]\colon a_{i,j}\mapsto a_{i,j},$ (13)
we have
@@
3.2.2 On augmentations of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$ for $m\geq 0$
Fix any $B\in B_{n}$ whose closure gives a knot $K$ in $\mathbb{R}^{3}$. We consider a braid $B^{\#}_{m}$ on $(n+2)$ strands defined by
$B^{\#}_{m}\coloneqq\sigma_{n}\cdot(\sigma_{n+1})^{2m+1}\cdot B\in B_{n+2}$
for every $m\in\mathbb{Z}_{\geq 0}$. Then, the closure of $B^{\#}_{m}$ gives the connected sum $K\#T_{(2,2m+1)}$ of $K$ and the $(2,2m+1)$-torus knot. See Figure 1 in the case of $m=1$. (When $m=0$, $T_{(2,1)}$ is the unknot.)